OpenStudy (anonymous):

a spherical balloon has a circumference of 21 cm. what is the approximate surface area of the balloon to the nearest square centimeter?

OpenStudy (campbell_st):

use the circumference of the sphere to find the radius \[21 = 2\times \pi \times r\] when you get a value substitute it into the surface area formula \[SA = 4 \pi r^2\] and alternate method to use \[SA = \frac{4 \pi^2 r^2}{\pi} \] which is \[SA = \frac{(2\pi r)^2}{\pi}\] or \[SA = \frac{C^2}{\pi}\]